OpenStudy (anonymous):

Tabitha wants to hang a painting in a gallery. The painting and frame must have an area of 58 square feet. The painting is 7 feet wide by 8 feet long. Which quadratic equation can be used to determine the thickness of the frame, x?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@TuringTest @dan815

OpenStudy (anonymous):

I'm not looking for a straight answer, I want to know how to do this.

OpenStudy (anonymous):

@amistre64

OpenStudy (mahmood786123):

7 times 8

OpenStudy (mahmood786123):

I think

OpenStudy (anonymous):

It can't be, that's the area of the painting itself.

OpenStudy (anonymous):

@AquaStone316

OpenStudy (anonymous):

I honestly don't know.

OpenStudy (welshfella):

( 7 +x)(8 + x) = 58

OpenStudy (welshfella):

ehere x is the thickness of the frame

OpenStudy (anonymous):

*facepalms* I forgot. Here are the answer choices x^2 + 15x - 2 = 0 x^2 + 15x + 58 = 0 4x^2 + 30x - 2 = 0 4x^2 + 30x + 58 = 0

OpenStudy (welshfella):

x^2 + 15x + 56 - 58 = 0

OpenStudy (welshfella):

do you see which one it is now?

OpenStudy (anonymous):

X^2+15x-2?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Thanks.